Home

The STAR experiment

  • Blogs
  • Events
  • Help&FAQ
  • Home
  • Polls
  • Sitemap
  • PWG
  • Software & Computing
  • Sub-systems
  • Council
  • General
  • Juniors

Detector Sub-systems

  • BEMC
    • BEMC Detector Operator Manual
    • Calibrations
      • BPRS
      • BSMD
      • BTOW - Calibration Procedure
      • Run 12 BTOW Calibration
      • Run 3 BTOW Calibration
      • Run 4 BTOW Calibration
      • Run 5 BTOW Calibration
      • Run 6 BTOW Calibration
      • Run 7 BTOW Calibration
      • Run 8 BTOW Calibration (2008)
      • Run 9 BTOW Calibration
        • 01 BTOW HV Mapping
        • 02 Comparing 2007, 2008 and 2009 BTOW Slopes
        • 03 study of 2009 slopes (jan)
        • 04 Spectra from Problem PMT Channels
        • 05 Summary of HV Adjustment Procedure
        • 06 comparison of BTOW status bits L2ped vs. L2W , pp 500 (jan)
        • 07 BTOW status tables ver 1, uploaded to DB, pp 500
        • 08 End or run status
        • 09 MIP peaks calculated using L2W stream
        • 10 Electron E/p from pp500 L2W events
        • 12 Correcting Relative gains from 500 GeV L2W
        • 13 Updating Calibration using the latest L2W production
        • 14 200 GeV Calibration
    • Database
    • Hardware
    • Mapping
    • Service Tasks
    • Software
    • Useful Documents
  • EEMC
  • ETOF
  • FCS
  • FGT
  • FPD & FMS & FPS
  • FTPC
  • FTT
  • HLT
  • L3
  • PMD
  • PP2PP
  • RICH
  • Roman Pot Phase II*
  • SSD
  • SVT
  • Slow Controls
  • TPC
  • TRG
  • Trigger Detectors
  • VPD
  • test

User login

Navigation

  • Groups
  • Recent posts
Home » Detector Sub-systems » BEMC » Calibrations

Run 9 BTOW Calibration

Updated on Tue, 2009-03-10 14:16. Originally created by mattheww on 2009-03-10 14:16. Under:
  • BEMC

Parent page for BTOW 2009 Calibrations

  • 01 BTOW HV Mapping
  • 02 Comparing 2007, 2008 and 2009 BTOW Slopes
  • 03 study of 2009 slopes (jan)
  • 04 Spectra from Problem PMT Channels
  • 05 Summary of HV Adjustment Procedure
  • 06 comparison of BTOW status bits L2ped vs. L2W , pp 500 (jan)
  • 07 BTOW status tables ver 1, uploaded to DB, pp 500
  • 08 End or run status
  • 09 MIP peaks calculated using L2W stream
  • 10 Electron E/p from pp500 L2W events
  • 12 Correcting Relative gains from 500 GeV L2W
  • 13 Updating Calibration using the latest L2W production
  • 14 200 GeV Calibration
‹ 08 Completed Calibration and Uncertainty up 01 BTOW HV Mapping ›
»
  • Printer-friendly version
  • Login or register to post comments
Privacy and Security notice webmaster